# Explicit quantum surrogates for quantum kernel models

### v1

@@ Appendix A Automatic Quantum Circuit Encoding
In this section, we briefly summarize the automatic quantum circuit encoding algorithm (AQCE) [30]. AQCE is an algorithm that produces a quantum circuit $\mathcal{C}$, which outputs a quantum state $\mathcal{C}|\bm{0}\rangle$ equivalent to a given arbitrary quantum state $|\Psi\rangle$ with the accuracy desired by the user. AQCE sequentially updates the unitary gates that construct $\mathcal{C}$ using a gradient-free method to maximize the fidelity function $F=\left|\left\langle 0\left|\hat{\mathcal{C}}^{\dagger}\right|\Psi\right\rangle\right|$.
@@
Here, $\bar{\mathbb{I}}_{m}$ is the complement of the subsystem $\mathbb{I}_{m}$ in the total qubit system. If we represent the fidelity tensor operator $\mathcal{F}_{m}$ and the unitary gate $\bm{U}_{m}$ in matrix form as $\bm{F}_{m}$ and $\bm{U}_{m}$, respectively, the expression is given by
$F_{m}=\left|\operatorname{Tr}\left[\bm{F}_{m}\bm{U}_{m}\right]\right|.$ (7)
@@
The pseudocode for the AQCE algorithm is presented in Alg. S1. While the process in line 7 is computationally expensive on classical computers, it can be efficiently computed using a quantum computer with the Hadamard test [30]. In contrast, the calculations from lines 8 to 10 can be efficiently performed on a classical computer.
@@
Inputs: Quantum state $|\Psi\rangle$, initial number of unitary gates $J_{0}\in\mathbb{N}$, increase in the number of unitary gates per sweep $\delta J\in\mathbb{N}$, number of sweep $N\in\mathbb{N}$, maximum number of unitary gates $J_{\rm max}\in\mathbb{N}$, target fidelity $F_{\rm target}\in(0,1]$, set of indices where a two-qubit unitary can be placed $\mathbb{B}$
@@
2:while $J<J_{\max}\wedge F<F_{\text{target}}$ do
@@
10:              Calculate $\bm{U}_{m}^{*}=\bm{X}_{l^{*}}\bm{Y}_{l^{*}}$ and determine the unitary gate $\mathcal{U}^{*}_{m,l^{*}}$ that corresponds to $\bm{U}^{*}_{m}$ and acts on $\mathbb{I}_{l}$.
@@
Appendix B Extending AQCE to isometries
@@
Using $\mathbb{I}_{m}$ defined in Appendix A, $F_{m}$ can be rewritten as
@@
$F_{m}=\sum_{k=0}^{K-1}\left|\operatorname{Tr}\left[\bm{F}_{m}^{(k)}\bm{U}_{m}\right]\right|.$ (19)
To optimize the unitary gate $\mathcal{U}_{m}$ to maximize the fidelity function $F_{m}$ using the AQCE algorithm strategy, we reformulate the fidelity function $F_{m}$ in the same form as in Eq. (7). This can be achieved if $\operatorname{Tr}\left[\bm{F}_{m}^{(k)}\bm{U}_{m}\right]$ is always real and positive in all cases. To achieve this, we transform $|\Psi^{(k)}\rangle$ as follows to cancel its global phase:
@@
$F_{m}$ $=\sum_{k=0}^{K-1}\left|\operatorname{Tr}\left[e^{i\theta_{m}^{(k)}}\bm{F}_{m}^{(k)}\bm{U}_{m}\right]\right|=\left|\operatorname{Tr}\left[\bm{F}_{m}\bm{U}_{m}\right]\right|,$ (21)
@@
1:Inputs: Set of orthogonal quantum states $\{|\Psi^{(k)}\rangle\}_{k}$, initial number of unitary gates $J_{0}\in\mathbb{N}$, increase in the number of unitary gates per sweep $\delta J\in\mathbb{N}$, number of sweep $N\in\mathbb{N}$, maximum number of unitary gates $J_{\rm max}\in\mathbb{N}$, set of target fidelities $\left\{F_{\rm target}^{(k)}\right\}_{k}$, $F_{\rm target}^{(k)}\in(0,1]$ for the quantum state $|\Psi^{(k)}\rangle$, set of indices where a two-qubit unitary can be placed $\mathbb{B}$
@@
4:while $J<J_{\max}\wedge F^{(0)}<F_{\text{target }}^{(0)}\wedge\cdots\wedge F^{(K-1)}<F_{\text{target }}^{(K-1)}$ do
@@
6:     $\mathcal{J}\leftarrow J+\delta J$
7:     for $\rm counter=1$ to $N$ do
@@
Appendix C Detailed conditions of numerical experiments
First, we provide the details of the numerical experiments in Fig. 2. As our implicit model, we employ a kernel support vector machine (SVM) model [35, 36]. We employ the one-vs-rest approach for multi-class classification; that is, we train a separate SVM model $f_{\mathrm{implicit}}^{(l)}$ for each class $l$ to distinguish it from all other classes. The overall prediction for an input $\bm{x}$ is made by outputting $\mathrm{argmax}_{l}f_{\mathrm{implicit}}^{(l)}(\bm{x})$. The SVM model is trained using scikit-learn 1.3.0 [42] with a regularization strength of $C=1.0$. To construct EQS, we first perform an exact diagonalization of $O_{\bm{\alpha},\mathcal{D}}^{(l)}$ to determine its eigenvalues $\lambda^{(l)}$ and eigenvectors |$\ket{\lambda^{(l)}}$. Then, we generate quantum circuits $\mathcal{C}^{(l)}$ using Alg. S2. The parameters for the algorithms are set as $F_{\text{target }}^{(0)}=0.6,\cdots,F_{\text{target }}^{(K-1)}=0.6$, $J_{0}=12$ and $\delta J=6$, without specifying $J_{\rm max}$. We run noiseless quantum circuit simulations using Qulacs 0.5.6 [41]. For the input data, we focus on the MNIST dataset from the MNISQ dataset [31], which has a fidelity of over 95%. We sample 1,000 data points for each label from the dataset, totaling 10,000 data points across all labels. Half of the data from each label is allocated for training, with the remaining half designated for testing. Additionally, we focus on 12-qubit VQE-generated datasets [14]. The datasets comprise six labels. For each label, we use all 300 data points, resulting in a total of 1,800 data points. Half of the data for each label is allocated for training, with the remaining half used for testing.
Next, we provide the details for the numerical experiments in Fig. 3. In these experiments, we focus on the MNISQ-MNIST dataset [31] with fidelity condition 95%. Let $\mathcal{C}^{(l)}(\bm{\theta})$ be a parameterized quantum circuit obtained by treating the two-qubit unitary gates in $\mathcal{C}^{(l)}$ of EQS as parameterized arbitrary two-qubit unitary gates with 15 independent parameters. The number of parameters is approximately $2,000$ to $3,000$, depending on the structure found via AQCE. Furthermore, let
@@ Appendix C Detailed conditions of numerical experiments
where $O^{(l)}=\sum_{k=0}^{K-1}\lambda_{k}^{(l)}|k\rangle\langle k|$, and
$p^{(l)}(\bm{x};\bm{\theta})=\frac{1}{1+e^{-f^{(l)}(\bm{x}_{m};\bm{\theta})}}.$ (23)
We define the weighted cross-entropy function as the loss on an input dataset $\mathcal{X}$:
@@
Here, $M=|\mathcal{X}|$, $M_{l}$ is the number of data points belonging to the label $l$, and $M_{\neq l}$ is the number of data points belonging to other labels, that is, $M_{\neq l}=\sum_{l^{\prime}\neq l}M_{l^{\prime}}$. $y_{\bm{x}}\in\{0,1\}$ is the label corresponding to the input data $\bm{x}$, where data belonging to $l$ are labeled as $1$, and all others are labeled as 00. The gradients of the loss function, $\frac{\partial L^{(l)}}{\partial\bm{\theta}}$, are computed using Qulacs 0.5.6 [41] and JAX 0.4.30 [43]. For $\mathcal{X}$, we sample $M_{l}=500$ data points for each label from the dataset that were not used in the construction of EQS, totaling $M=5000$ data points across all labels. The quantum circuit simulation is performed under a noiseless environment. Half of the data from each label is used for evaluating the gradients.
Appendix D Additional numerical experiments
D.1 Experiments on additional datasets
Here, we present the results for the MNISQ-Fashion MNIST dataset [31] and an 8-qubit VQE-generated dataset [14]. The results are shown in Fig. S1. The simulation conditions are identical to those described in Appendix C. In Fig. S1 (a), we can see that the accuracy of the exact low-rank model improves and approaches the accuracy of the original implicit model as $K$ increases, which is consistent with the results discussed in the main text. For example, the exact low-rank model exhibits only a 0.004 decrease in accuracy when compared to the implicit model at $K=10$ for the MNISQ-Fashion MNIST dataset. For the VQE-generated dataset in Fig. S1 (b), the accuracy of the exact low-rank model saturates already at $K=1$. This may be because the accuracy is 0.998 at $K=1$, sufficiently close to 1.0, leaving no room for improvement in accuracy with the increase of $K$. The reason for such a high accuracy may be attributed to factors discussed in the main text; in the case of the VQE-generated dataset, the first eigenvector of the observable in the implicit model exhibits high fidelity corresponding to its label.

### v2

@@ Appendix F Automatic quantum circuit encoding
In this section, we briefly summarize the automatic quantum circuit encoding algorithm (AQCE) [Shirakawa2021-uj]. AQCE is an algorithm that produces a quantum circuit $\mathcal{C}$, which outputs a quantum state $\mathcal{C}|0\rangle$ equivalent to a given arbitrary quantum state $|\Psi\rangle$ with the accuracy desired by the user. AQCE sequentially updates the unitary gates that construct $\mathcal{C}$ using a gradient-free method to maximize the fidelity function $F=\left|\left\langle 0\left|\mathcal{C}^{\dagger}\right|\Psi\right\rangle\right|$.
@@
Here, $\bar{\mathbb{I}}_{m}$ is the complement of the subsystem $\mathbb{I}_{m}$ in the total qubit system. If we represent the fidelity tensor operator $\mathcal{F}_{m}$ and the unitary gate $\mathcal{U}_{m}$ in matrix form as $\bm{F}_{m}$ and $\bm{U}_{m}$, respectively, the expression is given by
$F_{m}=\left|\operatorname{Tr}\left[\bm{F}_{m}\bm{U}_{m}^{\dagger}\right]\right|.$ (10)
@@
The pseudocode for the AQCE algorithm is presented in Alg. S1. While the process in line 7 is computationally expensive on classical computers, it can be efficiently computed using a quantum computer with the Hadamard test [Shirakawa2021-uj]. In contrast, the calculations from lines 8 to 10 can be efficiently performed on a classical computer.
@@
Inputs: Quantum state $|\Psi\rangle$, initial number of unitary gates $J_{0}\in\mathbb{N}$, increase in the number of unitary gates per sweep $\delta J\in\mathbb{N}$, number of sweep $N\in\mathbb{N}$, maximum number of unitary gates $J_{\max}\in\mathbb{N}$, target fidelity $F_{\mathrm{target}}\in(0,1]$, set of indices where a two-qubit unitary can be placed $\mathbb{B}$
@@
2:while $J<J_{\max}\wedge F<F_{\mathrm{target}}$ do
@@
10:     Calculate $\bm{U}_{m}^{*}=\bm{X}_{l^{*}}\bm{Y}_{l^{*}}$ and determine the unitary gate $\mathcal{U}^{*}_{m,l^{*}}$ that corresponds to $\bm{U}^{*}_{m}$ and acts on $\mathbb{I}_{l^{*}}$.
@@
Appendix G Extending AQCE to isometries
@@
Using $\mathbb{I}_{m}$ defined in Appendix F, $F_{m}$ can be rewritten as
@@
$F_{m}=\sum_{k=0}^{K-1}\left|\operatorname{Tr}\left[\bm{F}_{m}^{(k)}\bm{U}_{m}^{\dagger}\right]\right|.$ (22)
To optimize the unitary gate $\mathcal{U}_{m}$ to maximize the fidelity function $F_{m}$ using the AQCE algorithm strategy, we reformulate the fidelity function $F_{m}$ in the same form as in Eq. (10). This can be achieved if $\operatorname{Tr}\left[\bm{F}_{m}^{(k)}\bm{U}_{m}^{\dagger}\right]$ is always real and positive in all cases. To achieve this, we transform $|\Psi^{(k)}\rangle$ as follows to cancel its global phase:
@@
$F_{m}$ $=\sum_{k=0}^{K-1}\left|\operatorname{Tr}\left[e^{i\theta_{m}^{(k)}}\bm{F}_{m}^{(k)}\bm{U}_{m}^{\dagger}\right]\right|=\left|\operatorname{Tr}\left[\bm{F}_{m}\bm{U}_{m}^{\dagger}\right]\right|,$ (24)
@@
1:Inputs: Set of orthogonal quantum states $\{|\Psi^{(k)}\rangle\}_{k}$, initial number of unitary gates $J_{0}\in\mathbb{N}$, increase in the number of unitary gates per sweep $\delta J\in\mathbb{N}$, number of sweep $N\in\mathbb{N}$, maximum number of unitary gates $J_{\max}\in\mathbb{N}$, set of target fidelities $\left\{F_{\mathrm{target}}^{(k)}\right\}_{k}$, $F_{\mathrm{target}}^{(k)}\in(0,1]$ for the quantum state $|\Psi^{(k)}\rangle$, set of indices where a two-qubit unitary can be placed $\mathbb{B}$
@@
4:while $J<J_{\max}\ \wedge\ \min_{k}\,F^{(k)}<F^{(k)}_{\mathrm{target}}$ do
@@
6:  $J\leftarrow J+\delta J$
7:  for $\mathrm{counter}=1$ to NN do
@@
Appendix H Detailed conditions of numerical experiments
H.1 Dataset descriptions
Here we provide further details on the datasets used in our numerical experiments. For the tasks we consider, each input data point $\bm{x}$ is itself a quantum circuit, provided as a character string in QASM format [Cross2022-pb]. Consequently, the feature encoding map $U(\bm{x})$ is simply the execution of the circuit defined by that string. We selected two datasets with distinct structural properties to test our framework under different conditions.
H.1.1 MNISQ dataset
The MNISQ dataset [Placidi2023-kq] consists of 10-qubit quantum circuits that encode classical image vectors from the MNIST handwritten digit dataset [Lecun1998-dr]. The circuits were originally generated using the automatic quantum circuit encoding (AQCE) algorithm [Shirakawa2021-uj], which constructed each circuit $U$ such that its output state $U|0\rangle\langle 0|U^{\dagger}$ represents the corresponding classical image vector. A key property of this dataset is that circuits with the same digit label are designed to produce similar quantum states, representing a realistic scenario where intra-class similarity is high.
H.1.2 VQE-generated dataset
The VQE-generated dataset [Nakayama2023-zc] is composed of 4- to 20-qubit quantum circuits obtained from executing the variational quantum eigensolver (VQE) algorithm [Peruzzo2014-gz, Tilly2022-nh]. In contrast to the MNISQ dataset, this dataset is derived from a quantum-native problem: the classification of physically meaningful quantum states. Each circuit prepares an approximate ground state for a specific physical Hamiltonian (e.g., the transverse-field Ising model). A defining characteristic of this dataset is that states corresponding to different classes (i.e., ground states of different hamiltonians) are known to be nearly orthogonal. This provides a distinct test case for our framework on a quantum-native classification task.
H.2 SVM and EQS construction details
[…]
Next, building on the base EQS models constructed as described above, we now provide the specific details for the numerical experiments in Fig. 3. These experiments focus on the VQE-generated datasets to perform a scaling analysis with the number of qubits $n$. The core of our experiment is to compare the initial loss gradients under different initialization schemes for each target label $l$. To do this, we first take the circuit ansatz $\mathcal{C}^{(l)}$ found by our extended AQCE algorithm. We then treat it as a parameterized quantum circuit (PQC), denoted $\mathcal{C}^{(l)}(\bm{\theta})$, by considering each of its two-qubit unitary gates as an arbitrary two-qubit unitary with 15 independent parameters. We then precisely identify the advantage of our strategy by comparing the gradients under three distinct initialization schemes for the parameters $\bm{\theta}$:
  1. 1.
    EQS-initialized model: This is our proposed method. We set the parameters $\bm{\theta}$ to the specific values originally determined by our extended AQCE algorithm.
  2. 2.
    EQS ansatz with random parameters: As a control to test the benefit of the circuit structure alone, we use the same PQC ansatz $\mathcal{C}^{(l)}(\bm{\theta})$ but initialize its parameters $\bm{\theta}$ randomly from the interval $0,2\pi)$.
  3. 3.
    Random ansatz with random parameters: As a baseline, we construct a PQC ansatz of the same gate count as that of $\mathcal{C}^{(l)}(\bm{\theta})$, but with randomly chosen gate positions. Its parameters are also initialized randomly from $0,2\pi)$.
To calculate the gradients for these models, we first define the components leading to our loss function. First, the raw output of the model for label $l$ is given by:
@@ H.2 SVM and EQS construction details
where $O^{(l)}=\sum_{k=0}^{K-1}\lambda_{k}^{(l)}|k\rangle\langle k|$. Next, this output is mapped to the interval $(0,1)$ using the sigmoid function, yielding a value $p^{(l)}$:
$p^{(l)}(\bm{x};\bm{\theta})=\frac{1}{1+e^{-f^{(l)}(\bm{x};\bm{\theta})}}.$ (26)
Finally, using this value, we define the weighted cross-entropy function as the loss on an input dataset $\mathcal{X}$:
@@
Here, $M=|\mathcal{X}|$, $M_{l}$ is the number of data points belonging to the label $l$, and $M_{\neq l}$ is the number of data points belonging to other labels, that is, $M_{\neq l}=\sum_{l^{\prime}\neq l}M_{l^{\prime}}$. $y_{\bm{x}}\in\{0,1\}$ is the label corresponding to the input data $\bm{x}$, where data belonging to $l$ are labeled as $1$, and all others are labeled as 0. The gradients of the loss function, $\frac{\partial L^{(l)}}{\partial\bm{\theta}}$, are computed using Qulacs 0.5.6 [Suzuki2021-nj] and JAX 0.4.30 [jax2018github]. For $\mathcal{X}$, we sample $M_{l}=150$ data points for each label from the dataset that were not used in the training of EQS, resulting in a total of $M=750$ data points for the 5-label datasets (4-qubit) and $M=900$ for the 6-label dataset (8-, 12-, and 16-qubit). The quantum circuit simulation is performed under a noiseless environment.
Appendix I Experiments on additional datasets
To demonstrate the generality of our findings in Fig. 2, we repeated the same performance analysis on two additional datasets: the MNISQ-Fashion MNIST dataset [Placidi2023-kq] and an 8-qubit VQE-generated dataset [Nakayama2023-zc]. The simulation conditions were identical to those described in Appendix H.2, with only the datasets being replaced.
The results of these experiments, presented in Fig. S3, are consistent with those in the main text. For the MNISQ-Fashion MNIST dataset (Fig. S3 (a)), the accuracy of the exact low-rank model improves and approaches the accuracy of the original implicit model as $K$ increases, showing only a 0.004 decrease compared to the full model at $K=10$. In the case of the 8-qubit VQE-generated dataset (Fig. S3 (b)), the accuracy of the exact low-rank model saturates already at $K=1$. This high accuracy is attributed to the unique structure of the VQE-generated dataset, as discussed in the main text. Since states with different labels are nearly orthogonal, classification can be effectively achieved using only the first eigenvector, which has high fidelity with states of the corresponding label, as shown in Fig. S6 in Appendix K.
Furthermore, to confirm that the barren plateau mitigation effect shown in Fig. 3 is also a general phenomenon, we performed a gradient analysis on the 10-qubit MNISQ-MNIST dataset, following the same experimental protocol detailed in Appendix H.2. As presented in Fig. S4, the results show a multi-order-of-magnitude gradient enhancement for the EQS-initialized model compared to randomly initialized models. This is consistent with the findings for the VQE-generated datasets [Nakayama2023-zc] discussed in the main text, and confirms that our initialization strategy may be effective across different data environments.